[Brahms-l] Shift report 20040222 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 22 2004 - 07:54:00 EST
------------------------------------------------------------
Date: Sun Feb 22 07:53:59 2004, Shift: 00:00-08:00
Supervisor: DO, with: 
Log book#15 pages 64 - 64

Shift Summary:
------------------------------------------------------------
1. End of store from previous shift at 1am, 2nd store at 2am
   until 6:30am.

2. Data:
   MRS at 40d B1050: 680k tr3 runs 10479 -> 10488
   FS  at  4d B1/1 : 180k tr2 runs 10479 -> 10483 
           4d B1/2 : 437k tr2 runs 10484 -> 10488

   tpc ped: 10480, 10489
   others : 10481  10490

3. No detector problems. I just noticed that the event rate
   shown by the DAQ control panel is kinda oscillating 
   from expected average (100ev/s) to really low values, 
   even a few times as low as 0ev/s (DAQ in a "paused" state).
   This happens every 10-15s or so.    
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 22 07:56:53 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 22 2004 - 07:57:15 EST